OpenStudy (anonymous):

Erica is in training for a cross country race. She has been training by running 3 miles each day. She will be increasing her running by this weekly formula: y = 3(1.1)x. What is the value if x = 0? A. 0 B. 1.1 C. 3 D. 3.3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

y = 1.1 x + 3 y = 1.1 (0) + 3 y = 3